Calculate the average kinetic energy translation of an ideal gas at 270C
Answer #1

The average kinetic energy translation of an ideal gas at

Temperature T = 270 c

We need to convert temperature degree Celsius to kelvin

Therfore T = (270 + 273)

T = 543 Kelvin

Now the kinetic energy of an ideal gas can be written as

Where kb = boltzman constant

kb = 1.38 *10^-23 J/K

Now kinetic energy can become

K.E = 1124.01 *10^-23 J

This is the average kinetic energy of the ideal gas at 270 C